# Statistical physics analysis of graph neural networks: Approaching optimality in the contextual stochastic block model

### v2

@@ A.2 Self-consistent equations
We introduced the covariance $\operatorname{Cov}_{P}(x)=\mathbb{E}_{P}(xx^{T})-\mathbb{E}_{P}(x)\mathbb{E}_{P}(x^{T})$ and the tensorial product $x^{\otimes 2}=xx^{T}$. We used Stein’s lemma to simplify the differentials of $Q^{1/2}$ and $\hat{Q}^{1/2}$ and to transform the expression of $\hat{V}_{w}$ into a more accurate expression for numerical computation in terms of covariance. We used the identities $2x^{T}Q^{1/2}\frac{\partial Q^{1/2}}{\partial E_{k,l}}x=x^{T}E_{k,l}x$ and $-x^{T}V\frac{\partial V^{-1}}{\partial E_{k,l}}Vx=x^{T}E_{k,l}x$ for any element matrix $E_{k,l}$ and for any vector $x$. We have also that $\nabla_{V}\log\det V=V^{-1}$, considering its comatrix. Last we kept the first order in $\beta$ with the approximations $Q+V\approx Q$ and $\hat{Q}-\hat{V}\approx\hat{Q}$.
@@ B.1 Derivation of the free energy
The free energy is $-\beta Nf=\partial_{n}\mathbb{E}_{u,\Xi^{\mathrm{e}},W,y}Z^{n}(n=0)$ where the partition function is
@@
$Q_{q,kl}^{ab}=Q_{q,kl}+V_{q,kl}\delta_{a,b}\ ,$ $\hat{Q}_{q,kk}^{ab}=-\hat{Q}_{q,kk}+\frac{1}{2}(\hat{V}_{q,kk}+\hat{Q}_{q,kk})\delta_{a,b}\ ,$ $\hat{Q}_{q,kk}^{ab}=-\hat{Q}_{q,kk}+\hat{V}_{q,kk}\delta_{a,b}\ ,$ (210)
@@
::=:= $\ \int\mathrm{d}\Theta\mathrm{d}\hat{\Theta}e^{N\phi^{(n)}(\Theta,\hat{\Theta})}\ ,$ (215)
@@
$=$ $\operatorname*{extr}_{\Theta,\hat{\Theta}}\frac{\partial}{\partial n}(n=0)\phi^{(n)}(\Theta,\hat{\Theta})$ (219)
::=:= $\operatorname*{extr}_{\Theta,\hat{\Theta}}\phi(\Theta,\hat{\Theta})\ ,$ (220)
@@
ϕ=italic-$\phi=$ $\frac{1}{2}(V_{w}\hat{V}_{w}+\hat{V}_{w}Q_{w}-V_{w}\hat{Q}_{w})+\frac{1}{2}\operatorname{Tr}(V_{q}\hat{V}_{q}+\hat{V}_{q}Q_{q}-V_{q}\hat{Q}_{q})+\frac{1}{2}\operatorname{Tr}(V_{h}\hat{V}_{h}+\hat{V}_{h}Q_{h}-V_{h}\hat{Q}_{h})$ (221)
@@
We assume that $\ell$ is quadratic so $\psi_{qh}$ can be written as the following quadratic potential. Later we will take the limit $r\to\infty$ where $h$ is small and where $\ell$ can effectively be expanded around 00 as a quadratic potential.
@@
$\hat{m}\to\beta\hat{m}\ ,$ $\hat{Q}_{h}\to\beta^{2}\hat{Q}_{h}\ ,$ $\hat{V}_{h}\to\beta\hat{V}_{h}\ ,$ (239)
@@ B.2 Self-consistent equations
($\ -\mathrm{i}V_{qh}&\cdot\end{smallmatrix}\right)=t\mathcal{P}\left(G_{0}^{-1}D_{qh}G^{-1}\right)$ (251)
@@ B.3 Solution in the continuous limit at large $r$
$\varphi(x)$ =$=\left\{\begin{array}[]{cc}e^{\lambda^{\mathrm{e}}tx}&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \sum_{\nu>0}^{\infty}\nu(\lambda^{\mathrm{e}})^{\nu-1}\frac{I_{\nu}(2tx)}{tx}&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.\ ,$ (256)
$\Phi(x,z)$ =$=\left\{\begin{array}[]{cc}I_{0}(2t\sqrt{xz})&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \frac{I_{1}(2t(x+z))}{t(x+z)}&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.\ ,$ (259)
@@
$\qquad\left.{}+\int_{0,0^{-}}^{1^{+},1}\mathrm{d}x^{\prime\prime}\mathrm{d}z^{\prime\prime}\,\left(\hat{Q}_{h}(1-x^{\prime},x^{\prime\prime})+\mathcal{P}(\hat{m}^{\otimes 2})(1-x^{\prime},x^{\prime\prime})\right)D_{qh}^{-1}(x^{\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})\right]$
@@
$\qquad{}+\int_{0^{-},0}^{1,1^{+}}\mathrm{d}x^{\prime\prime}\mathrm{d}x^{\prime\prime\prime}\,G_{0}(x^{\prime},x^{\prime\prime})D_{qh}^{-1,T}(x^{\prime\prime},x^{\prime\prime\prime})\left(t\delta_{\mathrm{e}}Q_{qh}(x^{\prime\prime\prime},z^{\prime})+\mathcal{P}(\hat{m})(x^{\prime\prime\prime})(\lambda^{\mathrm{e}}tm(z^{\prime})+\sqrt{\mu}m_{w}\delta(z^{\prime}))\right)$
$\qquad{}+\int_{0,0^{-}}^{1^{+},1}\mathrm{d}z^{\prime\prime}\mathrm{d}z^{\prime\prime\prime}\,\left(t\delta_{\mathrm{e}}Q_{qh}(z^{\prime\prime\prime},x^{\prime})+(\lambda^{\mathrm{e}}tm(x^{\prime})+\sqrt{\mu}m_{w}\delta(x^{\prime}))\mathcal{P}(\hat{m})(z^{\prime\prime\prime})\right)D_{qh}^{-1}(z^{\prime\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})$
$\quad\left.{}+\int_{0^{-},0,0,0^{-}}^{1,1^{+},1^{+},1}\mathrm{d}x^{\prime\prime}\mathrm{d}x^{\prime\prime\prime}\mathrm{d}z^{\prime\prime}\mathrm{d}z^{\prime\prime\prime}\,G_{0}(x^{\prime},x^{\prime\prime})D_{qh}^{-1,T}(x^{\prime\prime},x^{\prime\prime\prime})\left(\hat{Q}_{h}(x^{\prime\prime\prime},z^{\prime\prime\prime})+\mathcal{P}(\hat{m}^{\otimes 2})(x^{\prime\prime\prime},z^{\prime\prime\prime})\right)D_{qh}^{-1}(z^{\prime\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})\right]\ ;$
@@ B.4 Higher orders in $1/r$: how to pursue the computation
where we remind that $G_{0}=t^{2}V_{h}+V_{w}\delta(x,z)=\mathcal{O}(1/r)$, $G_{h}=\hat{V}_{h}+\bar{s}\delta(1-x,1-z)$ and $D_{qh}=D-t\delta_{\mathrm{e}}V_{qh}^{T}$. These equations form a system of non-linear integral equations. A perturbative approach with expansion in powers of $1/r$ should allow to solve it. At each order one has to solve linear integral equations whose resolvant is $\Phi$ for $V_{h}$ and $\hat{V}_{h}$, the previously determined resolvant to the constant order. The perturbations have to summed and the resulting $V_{qh}$, $V_{h}$ and $\hat{V}_{h}$ can be used to express $h^{*}$, $h^{{}^{\prime}*}$ and the other order parameters.
@@ B.5 Interpretation of terms of DMFT: computation
$\quad=\frac{V_{w}}{N}\left\{\begin{array}[]{lc}\operatorname{Tr}\left(e^{\frac{tx}{\sqrt{N}}\tilde{A}}e^{\frac{tz}{\sqrt{N}}\tilde{A}^{T}}\right)&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \operatorname{Tr}\left(e^{\frac{tx+tz}{\sqrt{N}}\tilde{A}^{\mathrm{s}}}\right)&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.\ .$ (297)
@@
=$\quad=\left\{\begin{array}[]{lc}\theta(x-z)&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \theta(x-z)(x-z)^{-1}I_{1}(2t(x-z))&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.$ (309)

### v3

@@ A.2 Self-consistent equations
We introduced the covariance $\operatorname{Cov}_{P}(x)=\mathbb{E}_{P}(xx^{T})-\mathbb{E}_{P}(x)\mathbb{E}_{P}(x^{T})$ and the tensorial product $x^{\otimes 2}=xx^{T}$. We used Stein’s lemma to simplify the differentials of $Q^{1/2}$ and $\hat{Q}^{1/2}$ and to transform the expression of $\hat{V}_{w}$ into a more accurate expression for numerical computation in terms of covariance. We used the identities $2x^{T}Q^{1/2}\frac{\partial Q^{1/2}}{\partial E_{k,l}}x=x^{T}E_{k,l}x$ and $x^{T}V\frac{\partial V^{-1}}{\partial E_{k,l}}Vx=x^{T}E_{k,l}x$ for any element matrix $E_{k,l}$ and for any vector $x$. We have also that $\nabla_{V}\log\det V=V^{-1}$, considering its comatrix. Last we kept the first order in $\beta$ with the approximations $Q+V\approx Q$ and $\hat{Q}-\hat{V}\approx\hat{Q}$.
@@ B.1 Derivation of the free energy
The free energy is $\beta Nf=\partial_{n}\mathbb{E}_{u,\Xi^{\mathrm{e}},W,y}Z^{n}(n=0)$ where the partition function is
@@
$Q_{q,kl}^{ab}=Q_{q,kl}+V_{q,kl}\delta_{a,b}\ ,$ $\hat{Q}_{q,kk}^{ab}=-\hat{Q}_{q,kk}+\frac{1}{2}(\hat{V}_{q,kk}+\hat{Q}_{q,kk})\delta_{a,b}\ ,$ $\hat{Q}_{q,kl}^{ab}=-\hat{Q}_{q,kl}+\hat{V}_{q,kl}\delta_{a,b}\ ,$ (210)
@@
$:=$ $\ \int\mathrm{d}\Theta\mathrm{d}\hat{\Theta}e^{N\phi^{(n)}(\Theta,\hat{\Theta})}\$, (215)
@@
[…]
@@
$\phi=$ $\frac{1}{2}(V_{w}\hat{V}_{w}+\hat{V}_{w}Q_{w}-V_{w}\hat{Q}_{w})+\frac{1}{2}\operatorname{Tr}(V_{q}\hat{V}_{q}+\hat{V}_{q}Q_{q}-V_{q}\hat{Q}_{q})+\frac{1}{2}\operatorname{Tr}(V_{h}\hat{V}_{h}+\hat{V}_{h}Q_{h}-V_{h}\hat{Q}_{h})$ (221)
@@
We assume that $\ell$ is quadratic so $\psi_{qh}$ can be written as the following quadratic potential. Later we will take the limit $r\to\infty$ where $h$ is small and where $\ell$ can effectively be expanded around 0 as a quadratic potential.
@@
$\hat{m}\to\beta\hat{m}\$, $\hat{Q}_{h}\to\beta^{2}\hat{Q}_{h}\ ,$ $\hat{V}_{h}\to\beta\hat{V}_{h}\ ,$ (239)
@@ B.2 Self-consistent equations
$\left(\begin{smallmatrix}\cdot&\cdot\\ -\mathrm{i}V_{qh}&\cdot\end{smallmatrix}\right)=t\mathcal{P}\left(G_{0}^{-1}D_{qh}G^{-1}\right)$ (251)
@@ B.3 Solution in the continuous limit at large $r$
$\varphi(x)$ $=\left\{\begin{array}[]{cc}e^{\lambda^{\mathrm{e}}tx}&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \sum_{\nu>0}^{\infty}\nu(\lambda^{\mathrm{e}})^{\nu-1}\frac{I_{\nu}(2tx)}{tx}&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right$.\ , (256)
$\Phi(x,z)$ $=\left\{\begin{array}[]{cc}I_{0}(2t\sqrt{xz})&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \frac{I_{1}(2t(x+z))}{t(x+z)}&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right$.\ , (259)
@@
$\hskip 18.49988pt\left.{}+\int_{0,0^{-}}^{1^{+},1}\mathrm{d}x^{\prime\prime}\mathrm{d}z^{\prime\prime}\,\left(\hat{Q}_{h}(1-x^{\prime},x^{\prime\prime})+\mathcal{P}(\hat{m}^{\otimes 2})(1-x^{\prime},x^{\prime\prime})\right)D_{qh}^{-1}(x^{\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})\right]$
@@
$\hskip 18.49988pt{}+\int_{0^{-},0}^{1,1^{+}}\mathrm{d}x^{\prime\prime}\mathrm{d}x^{\prime\prime\prime}\,G_{0}(x^{\prime},x^{\prime\prime})D_{qh}^{-1,T}(x^{\prime\prime},x^{\prime\prime\prime})\left(t\delta_{\mathrm{e}}Q_{qh}(x^{\prime\prime\prime},z^{\prime})+\mathcal{P}(\hat{m})(x^{\prime\prime\prime})(\lambda^{\mathrm{e}}tm(z^{\prime})+\sqrt{\mu}m_{w}\delta(z^{\prime}))\right)$
$\hskip 18.49988pt{}+\int_{0,0^{-}}^{1^{+},1}\mathrm{d}z^{\prime\prime\prime}\mathrm{d}z^{\prime\prime}\,\left(t\delta_{\mathrm{e}}Q_{qh}(z^{\prime\prime\prime},x^{\prime})+(\lambda^{\mathrm{e}}tm(x^{\prime})+\sqrt{\mu}m_{w}\delta(x^{\prime}))\mathcal{P}(\hat{m})(z^{\prime\prime\prime})\right)D_{qh}^{-1}(z^{\prime\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})$
$\qquad\left.{}+\int_{0^{-},0,0,0^{-}}^{1,1^{+},1^{+},1}\mathrm{d}x^{\prime\prime}\mathrm{d}x^{\prime\prime\prime}\mathrm{d}z^{\prime\prime\prime}\mathrm{d}z^{\prime\prime}\,G_{0}(x^{\prime},x^{\prime\prime})D_{qh}^{-1,T}(x^{\prime\prime},x^{\prime\prime\prime})\left(\hat{Q}_{h}(x^{\prime\prime\prime},z^{\prime\prime\prime})+\mathcal{P}(\hat{m}^{\otimes 2})(x^{\prime\prime\prime},z^{\prime\prime\prime})\right)D_{qh}^{-1}(z^{\prime\prime\prime},z^{\prime\prime})G_{0}(z^{\prime\prime},z^{\prime})\right]\ ;$
@@ B.4 Higher orders in $1/r$: how to pursue the computation
where we remind that $G_{0}=t^{2}V_{h}+V_{w}\delta(x,z)=\mathcal{O}(1/r)$, $G_{h}=\hat{V}_{h}+\bar{s}\delta(1-x,1-z)$ and $D_{qh}=D-t\delta_{\mathrm{e}}V_{qh}^{T}$. These equations form a system of non-linear integral equations. A perturbative approach with expansion in powers of $1/r$ should allow to solve it. At each order one has to solve linear integral equations whose resolvant is $\Phi$ for $V_{h}$ and $\hat{V}_{h}$, the previously determined resolvant to the constant order. The perturbations have to summed and the resulting $V_{qh}$, $V_{h}$ and $\hat{V}_{h}$ can be used to express $^{*}, hh^{{}^{\prime}*}$ and the other order parameters.
@@ B.5 Interpretation of terms of DMFT: computation
$\quad=\frac{V_{w}}{N}\left\{\begin{array}[]{lc}\operatorname{Tr}\left(e^{\frac{tx}{\sqrt{N}}\tilde{A}}e^{\frac{tz}{\sqrt{N}}\tilde{A}^{T}}\right)&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \operatorname{Tr}\left(e^{\frac{tx+tz}{\sqrt{N}}\tilde{A}^{\mathrm{s}}}\right)&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right.\$. (297)
@@
$\quad=\left\{\begin{array}[]{lc}\theta(x-z)&\mathrm{if}\,\delta_{\mathrm{e}}=0\\ \theta(x-z)(x-z)^{-1}I_{1}(2t(x-z))&\mathrm{if}\,\delta_{\mathrm{e}}=1\end{array}\right$. (309)